OpenStudy (anonymous):

.calculate the uncertainity in momentum of an electron if it is confined to linear region of 10-10m? q2.in the hydrogen atom the energy of an electron in 1st bohrs orbit is 13.12*105J mol-1 ? what is the energy required for its excitation to bohrs second orbit?

OpenStudy (chmvijay):

can you write the formula to find uncertainity in momnetum

OpenStudy (anonymous):

i donot know the formula

OpenStudy (chmvijay):

ΔX . ΔP = h/4 *pie do u know this

OpenStudy (anonymous):

yessssssss par value samj nhi arhi kya kismein put karooo

OpenStudy (chmvijay):

delta P kaa matalb uncertainty in momntum tumko ΔX = 1 x 10-10 meter. ΔP = ? find karna hai h constant kaa value malum hai naa

OpenStudy (anonymous):

6.626*10raised to the power-34js divided by 4 pie hoga na

OpenStudy (chmvijay):

h = 6.64 x 10-34 joule chaloo calculate karke deltap kaa vlue batoo muje

OpenStudy (anonymous):

5.27*10-26kgms-1

OpenStudy (chmvijay):

aur eka baar check karo 10 raised ????check it once again

OpenStudy (anonymous):

10 raiesed to the power -26 hai

OpenStudy (chmvijay):

5.27 kyaaa 52.7 deko eka baar

OpenStudy (anonymous):

5.27 hai

OpenStudy (chmvijay):

nahiiiiiiiiiiiiiii = 6.626 *10 -34 / 12.57 *10-10 =0.528 x 10-24 kg.m sec-1

OpenStudy (anonymous):

but answer yhi de rakha hai

OpenStudy (chmvijay):

but this is right answer 0.528 x 10-24 kg.m sec-1

OpenStudy (anonymous):

r u sureeeee

OpenStudy (chmvijay):

LOL iam sureeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeee 100%

OpenStudy (anonymous):

kkkkk now next question plzz

OpenStudy (chmvijay):

that also u need to noo one formula try to find it

OpenStudy (anonymous):

do u know the second question i type

OpenStudy (chmvijay):

yaaa i know

OpenStudy (anonymous):

tehn plzz tell solution of tha t also

OpenStudy (chmvijay):

E =13.12*105J mol-1 / n^2 put n=2 and find E

OpenStudy (anonymous):

9.84*10 raised topwer 5 ayega

OpenStudy (chmvijay):

nahi yaar n^2 boletoo 2^2 kitna huaa

OpenStudy (anonymous):

4

OpenStudy (anonymous):

thereeee or nottttttt kya ayega naswer batona

OpenStudy (chmvijay):

13.12 / 4 kitna hua

OpenStudy (anonymous):

3.28

OpenStudy (chmvijay):

haa bara bara hai 3.28 *10^5

OpenStudy (anonymous):

kyooooo

OpenStudy (anonymous):

bataooooo ageeeeee

OpenStudy (anonymous):

aap na plzzzzzzz mujhe jaldi se questions ke answer batdo i will be thnkful to u then

OpenStudy (anonymous):

aap na plzzzzzzz mujhe jaldi se questions ke answer batdo i will be thnkful to u then

OpenStudy (anonymous):

aap na plzzzzzzz mujhe jaldi se questions ke answer batdo i will be thnkful to u then

OpenStudy (chmvijay):

are already maine btayaaa hai naaa aur kya chahiyeee

OpenStudy (anonymous):

donoooooo answer galat haii yarrr asnmnt mein kuch aur hai answer

OpenStudy (chmvijay):

answer too batawoo muje

OpenStudy (anonymous):

how much water is formed by mixing 100g h2 and 100 g 02????????

OpenStudy (chmvijay):

LOL

OpenStudy (anonymous):

kyaaaaa huaaaaaa ab yeh nxt question hai

OpenStudy (anonymous):

whttttt happnd nowwww

OpenStudy (anonymous):

thereeeee or notttttttt

thomaster (thomaster):

It's like your talking in a different language lol :O

OpenStudy (preetha):

This is so funny! A chemistry explanation in hindi on a site that claims to be English only! Another #onlyonOpenStudy moment!

OpenStudy (chmvijay):

Is it ok to use Hindi or should we be strict to English!! because some students understand better in their own language !!

OpenStudy (nincompoop):

Engdi is okay to use.

mathslover (mathslover):

Haha, Nice tag @Preetha mam : #onlyonOpenStudy Moment! :)

OpenStudy (anonymous):

The second question can be determined by looking at the bohr equation and determining the bohr radius for n=2 orbit. I don't think the application of a ratios would work in this case, as if I'm not mistaken E depends on the square of n. Good luck